40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Oct 7 14:51:20 2019, aaron, Update, Electronics, WFS head RF measurements WFShead_response.pdf
    Reply  Mon Oct 7 19:50:33 2019, gautam, Update, IOO, IMC locking not working after this work IMCflaky.pngIMG_8015.JPG
       Reply  Tue Oct 8 10:29:19 2019, gautam, Update, IOO, MC Transmission scan MC2_transmission_scatter.pdftransmissionMaps.pdfcorrelStructure.pdf
       Reply  Tue Oct 8 16:54:56 2019, rana, Update, IOO, IMC locking not working after this work 
    Reply  Tue Oct 8 16:00:06 2019, aaron, Update, Electronics, WFS head RF measurements WFS_ACresponse.pdfWFS_DCresponse.pdf
       Reply  Wed Oct 9 12:15:05 2019, rana, Update, Electronics, WFS head RF measurements 
Message ID: 14945     Entry time: Mon Oct 7 14:51:20 2019     Reply to this: 14946   14951
Author: aaron 
Type: Update 
Category: Electronics 
Subject: WFS head RF measurements 

Mon Oct 7 14:51:53 2019. I closed the PSL shutter to measure the WFS head responsivity.

I made a thru calibration as in this elog, treating laser, reference PD, and WFS RF output as a three-port device. The DC current supplied to the laser is 20.0 mA in all cases. The Agilent spectrum analyzer supplies a -10 dBm excitation to Jenne laser's AM port, and A/B is measured with 20dB attenuation on each input port. Results are in /users/aaron/WFS/data/191007/. The calibration had 100 averages, all other measurements 32 averages; other parameters found in the yml file, same folder as the data.

Measurement Reference PD DC (V) WFS Segment DC (V) WFS Segment DC, beam blocked (V) File Notes
WFS 1 Segment 1 1.86 0.79 -0.23
TFAG4395A_07-10-2019_154446.txt
 
WFS 1 Segment 2 1.86 0.72 -0.30 TFAG4395A_07-10-2019_155017.txt  
WFS 1 Segment 3 1.86 0.79 -0.21
TFAG4395A_07-10-2019_155645.txt
 
WFS 1 Segment 4 1.86 0.70 -0.30
TFAG4395A_07-10-2019_160334.txt
TFAG4395A_07-10-2019_160847.txt
I noticed the BS-PRM illuminator was on, and turned it off for the second measurement
WFS 2 Segment 1 1.86 0.56 -0.38 TFAG4395A_07-10-2019_162533.txt  
WFS 2 Segment 2 1.86 0.71 -0.21
TFAG4395A_07-10-2019_163402.txt
 
WFS 2 Segment 3 1.86 0.68 -0.28 TFAG4395A_07-10-2019_164152.txt  
WFS 2 Segment 4 1.86 0.57 -0.42 TFAG4395A_07-10-2019_164745.txt  

 

I normalized the result by the difference between the dark and bright DC levels of each segment.

Mon Oct 7 17:29:58 2019 opened PSL shutter.

Attachment 1: WFShead_response.pdf  54 kB  Uploaded Tue Oct 8 19:09:30 2019  | Hide | Hide all
WFShead_response.pdf
ELOG V3.1.3-